OpenStudy (anonymous):

I will be your fan and give you a medal What is the reciprocal of 1/5 A.-5 B.-1/5 C.1/5 D.5

OpenStudy (mathstudent55):

The reciprocal of a fraction is obtained by simply flipping it over.

OpenStudy (zzr0ck3r):

the ricipical of \(\frac{a}{b}\) is \(\frac{b}{a}\) note: \(a=\frac{a}{1}\)

OpenStudy (anonymous):

And that's not in the answer choices and so I think it's D

OpenStudy (zzr0ck3r):

yes d

OpenStudy (zzr0ck3r):

5/1 = 5

OpenStudy (anonymous):

Thank you